OpenStudy (anonymous):

The sine of 38 degrees is equal to 0.616, rounded to the nearest thousandth. What is the measure of an angle whose cosine is also equal to 0.616? 38 42 52 128

OpenStudy (anonymous):

If:

OpenStudy (anonymous):

i just wanted to see if the answer i got was correct i got 42

OpenStudy (anonymous):

My calculator says cos(42)=0.7431, so that wouldn't be correct.

OpenStudy (ranga):

sine and cosine are "co" functions. sin(x) = cos(90 - x) cos(x) = sin(90 - x)

OpenStudy (anonymous):

In a right triangle, the sine of an angle is equal to the ratio of the opposite leg of the triangle to that angle divided by the hypotenuse. The cosine is the same, except it's the adjacent leg of the triangle. In my diagram is there an angle that has an adjacent side that corresponds to the opposing side of 38 degrees?

OpenStudy (ranga):

Same is true of other "co" functions too: sec(x) and cosec(x) tan(x) and cot(x).

OpenStudy (ranga):

sin(38) is same as cos(90-38)

OpenStudy (anonymous):

52

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

the answer is 52